# On the Robustness of Kernel Goodness-of-Fit Tests

### v4

@@ 5.3 A Kernel Exponential Family Model for Density Estimation
Our next example concerns density estimation using a kernel exponential family (KEF) model (Canu and Smola,, 2006). Given a reproducing kernel $k_{0}:\mathbb{R}^{d}\times\mathbb{R}^{d}\to\mathbb{R}$ and a reference density $p_{0}$ on $\mathbb{R}^{d}$, we have $p_{\eta}(\mathbf{x})\propto p_{0}(\mathbf{x})\exp(-f_{\eta}(\mathbf{x}))$, where $f_{\eta}$ lies in the RKHS $\mathcal{H}_{k_{0}}$ associated with $k_{0}$. We follow the setting in Matsubara et al., (2022) which uses a finite-basis approximation of the RKHS function so that $f_{\eta}(\mathbf{x})=\sum_{l=1}^{25}\eta_{l}\phi_{l}(\mathbf{x})$, where $\eta=(\eta_{1},\ldots,\eta_{25})^{\top}\in\mathbb{R}^{25}$ and $\phi_{l}(\mathbf{x})=(x^{l}/\sqrt{l!})\exp(-x^{2}/2)$ for $l=1,2,\ldots,25$.
@@
We are interested in testing the goodness-of-fit of the KEF model under data contamination. Suppose that the practitioners believe the observed data are subject to contamination, so that a robust estimator for $\eta$ is used. Robust parameter estimation for exponential family models was studied in Barp et al., (2019), which proposed a minimum-distance estimator using KSD with tilted kernels. With a suitable weighting function, the resulting estimator is bias-robust (Barp et al.,, 2019, Proposition 7) and thus not susceptible to contamination. However, how to test the goodness-of-fit of the estimated model in this setting remains an open question. The standard IMQ-KSD test is not suitable since it is not robust and can thus falsely reject a good model due to contamination. Using a tilted kernel is also not enough since it is not quantitatively robust and hence can suffer from uncontrolled Type-I error. We therefore propose to use our robust-KSD test for this task. One challenge with the KEF model is that its density does not have a closed form, and although sampling from $P$ is feasible in this low-dimensional example ($d=1$), it would become challenging in high dimensions, making MMD-based tests not well-suited for this model. Our robust-KSD test, however, does not suffer from this limitation.
@@
We use the data set as Matsubara et al., (2022); Key et al., (2025), which is a $1$-dimensional data set of $82$ galaxy velocities (Postman et al.,, 1986; Roeder,, 1990). To avoid using the same data for model training and testing, we randomly split the data into equal halves, each containing $n_{\mathrm{data}}=41$ data points. To each half, we then add $n_{\mathrm{ol}}$ independent realisations from $R=\mathcal{N}(z,0.1^{2})$ for some value of $z$ for the mean contamination, so that the resulting data set has a size of $n=n_{\mathrm{data}}+n_{\mathrm{ol}}$ and a contamination ratio of $\epsilon=n_{\mathrm{ol}}/(n_{\mathrm{data}}+n_{\mathrm{ol}})$. We then fit the KEF model to one half of the data, and perform the tests on the other half. Both the robust minimum KSD estimation and the robust-KSD test use the same tilted kernel $k(x,y)=w(x)h_{\Lambda}(x-y)w(y)$ with $w(x)=(1+x^{2})^{-1/2}$ and a sum of IMQ kernels $h_{\Lambda}(x-y)=\sum_{\lambda^{2}\in\Lambda}(1+|x-y|^{2}/(2\lambda^{2}))^{-1/2}$ where $\Lambda=\{0.6,1,1.2\}$. The choice of the weighting function follows Matsubara et al., (2022) and the use of a sum kernel is recommended in Key et al., (2025). Direct computation shows that the Stein kernel $u_{p}$ is bounded in this case. We choose $\epsilon_{0}=0.2$ when computing $\theta$ for R-KSD.
@@ 5.4 Limitations for Multimodal Models
We investigate the performance of the robust test when the model has multimodality and the data are drawn from the same model with misspecified mixing ratios, thus revealing a limitation of our robust-KSD test. It is well-known that the standard KSD test performs poorly with mixture models with well-separated modes (Liu et al.,, 2023). This is intuitively due to the myopia of the score function, making the test blind to disparities in the mixing ratios, and it is a general issue for KSD and other score-based discrepancies (Wenliang and Kanagawa,, 2020; Zhang et al.,, 2022). We demonstrate that our robust test can also suffer from this limitation, which is made even more prominent due to the relaxation of the point null to a KSD-ball. We consider a 2-dimensional mixture of Gaussian model $P=\sum_{j=1}^{5}\pi_{j}\mathcal{N}(\gamma\mu_{j},I_{2})$, where $\pi_{j}\propto u_{j}$ with $u_{1},\ldots,u_{5}$ drawn independently from $\mathrm{Uniform}(0,1)$, each mean $\mu_{j}\in\mathbb{R}^{d}$ has entries drawn independently from $\mathrm{Uniform}(-2,2)$, and $\gamma>0$ is a scaling factor controlling the mode separation. The data-generating distribution $Q$ has a different set of mixture ratios that are randomly generated. For R-KSD, the uncertainty radius is chosen to be $\theta=\epsilon_{0}\tau_{\infty}^{1/2}$ with different values of $\epsilon_{0}$, and we label the corresponding result as R-KSD-$\epsilon_{0}$.
@@ 6 Conclusion and Discussion
This paper studied the robustness of the kernel GOF test of Liu et al., (2016); Chwialkowski et al., (2016). We showed that existing KSD-based tests lack qualitative robustness when stationary kernels are used, but achieve it when employing tilted kernels inspired by robust estimation methods (Barp et al.,, 2019; Matsubara et al.,, 2022). Since qualitative robustness alone does not ensure calibration under fixed contamination levels, we proposed a novel robust KSD test that provably controls the Type-I error when the data-generating distribution lies in a KSD-ball around the reference model, while consistently rejecting alternatives outside of it. We then discussed how to select the radius of this ball under different contamination models, namely Huber’s contamination model and the density-band model. Empirical results on synthetic and real data sets demonstrate that our robust GOF test is well-calibrated and has non-trivial power. We conclude by discussing promising directions for future work.
Firstly, we established consistency of the proposed robust KSD test against all alternative distributions outside the null set, i.e., $H^{\mathrm{C}}_{1}:Q\not\in\mathcal{B}^{\mathrm{KSD}}(P;\theta)$. In the broader robust testing literature, however, it is also common to restrict the alternative set to a neighborhood centered at a specific alternative $P^{\prime}\neq P$ and study minimax properties against such hypothesis (Huber,, 1965; Levy,, 2008; Gül and Zoubir,, 2016; Sun and Zou,, 2023). Understanding whether our robust KSD test achieves minimax optimality in this setting is an interesting future work.
Secondly, as discussed in Section˜2.3, our proposed robust KSD test does not require the Stein kernel $u_{p}$ to be bounded. A bounded Stein kernel is convenient as the resulting KSD-ball $\mathcal{B}^{\mathrm{KSD}}(P;\theta)$ contains many contamination models of interest, such as Huber’s model with an arbitrary noise distribution. However, this could be too strict in some applications where outliers cannot take an arbitrary value, e.g., when the domain is bounded or the data acquisition procedure is highly regulated to prevent extreme outliers. In these cases, even with an unbounded kernel, the KSD-ball could still capture all contamination models of interest. Unbounded Stein kernels possess appealing theoretical properties such as convergence-control of moments (Kanagawa et al.,, 2022). Exploring whether such kernels can lead to more powerful robust-KSD tests is another worthwhile future research direction.
Acknowledgements
@@
The appendix is organized as follows.
Throughout the appendix, given i.i.d. random variables $\mathbf{X}_{1},\ldots,\mathbf{X}_{n}\sim Q\in\mathcal{P}(\mathbb{R}^{d})$ and an integrable function $f:(\mathbb{R}^{d})^{n}\to\mathbb{R}$, we will write for brevity $\mathbb{E}[f(\mathbf{X}_{1},\ldots,\mathbf{X}_{n})]=\int_{\mathbb{R}^{d}}\cdots\int_{\mathbb{R}^{d}}f(\mathbf{x}_{1},\ldots,\mathbf{x}_{n})Q(\mathop{}\!\mathrm{d}\mathbf{x}_{1})\cdots Q(\mathop{}\!\mathrm{d}\mathbf{x}_{n})$.
@@ Appendix A Proofs of Theoretical Results
This section holds proofs of the theoretical results.
@@ A.1 Proof of Theorem˜1 and Related Preliminary Results
Assume $\sigma_{\infty}^{2}\coloneqq\sup_{\mathbf{z}\in\mathbb{R}^{d}}\mathbb{E}_{\mathbf{X}\sim P}[u_{p}(\mathbf{X},\mathbf{z})^{2}]<\infty$ and suppose ˜1 holds. Then, for any $\delta>0$ and integers $0<m^{\prime}<n$, the following event holds with probability at least $1-\delta$:
@@
Suppose ˜1 holds. Assume (i) $\sigma_{\infty}^{2}\coloneqq\sup_{\mathbf{z}\in\mathbb{R}^{d}}\mathbb{E}_{\mathbf{X}\sim P}[u_{p}(\mathbf{X},\mathbf{z})^{2}]<\infty$ and (ii) $\xi_{P}^{2}\coloneqq\mathbb{E}_{\mathbf{X}\sim P}[u_{p}(\mathbf{X},\mathbf{X})^{2}]<\infty$. Then there exist constants $C,C_{P,1}$, and $C_{P,2}$ depending only on $P$ such that, for any $\delta>0$ and integers $0<m^{\prime}<n$, the following event holds with probability at least $1-\delta$:
@@
There exists a constant $C>0$ such that the following holds. Suppose ˜1 holds. Also assume the following integrability conditions
@@ A.1.1 Proof of Lemma˜7
The test statistic can be decomposed in three terms: a term with uncorrupted samples, a term with corrupted samples, and a term for interactions between these two groups:
@@
Bounding the second term in (10) requires bounding $S_{m,\mathbf{z}}$. Because by assumption $\mathbf{X}_{i}=\mathbf{X}_{i}^{\ast}\sim P$ are i.i.d. for $i=1,\ldots,m$, the term $S_{m,\mathbf{z}}$ is a sum of i.i.d. random variables $u_{p}(\mathbf{X}_{i}^{\ast},\mathbf{z})$. Moreover, $u_{p}(\mathbf{X}_{i}^{\ast},\mathbf{z})$ are zero-mean, because the proof of Gorham and Mackey, (2017, Proposition 1) shows that $\mathbb{E}_{\mathbf{X}^{\ast}\sim P}[u_{p}(\mathbf{X}^{\ast},\cdot)]=0$ whenever $\mathbb{E}_{\mathbf{X}^{\ast}\sim P}[\|\mathbf{s}_{p}(\mathbf{X}^{\ast})\|_{2}]<\infty$, which holds under ˜1. We can therefore compute the variance of $S_{m,\mathbf{z}}$ as
@@ A.1.2 Proof of Lemma˜8
To prove this lemma, we will need the following concentration bound for quadratic forms of multinomial random vectors. This is a special instance of the Hanson-Wright inequality due to Adamczak, (2015) when the dependent random vectors follow multinomial distributions.
Lemma 10 (Hanson-Wright inequality; Adamczak, 2015, Theorem 2.5).
@@
where $\mu\coloneqq-n^{-1}\sum_{1\leq i,j\leq n}a_{ij}+\sum_{1\leq i\leq n}a_{ii}$, and $\|A\|_{\mathrm{F}}=(\sum_{1\leq i,j\leq n}a_{ij}^{2})^{1/2}$ is the Frobenius norm.
Proof of Lemma˜10  Adamczak, (2015, Remark 2.2) shows that multinomial random vectors satisfy the convex concentration inequality (Adamczak,, 2015, Definition 2.2). Therefore, we can apply Adamczak, (2015, Theorem 2.5) to conclude that there exists constants $C_{1},C_{2}$ such that, for any $n\times n$ matrix $A$ and any $t>0$,
@@
where $\|A\|_{\mathrm{op}}=\sup_{\|\mathbf{x}\|_{2}\leq 1}\|A\mathbf{x}\|_{2}$ is the operator norm, $\mu=\mathbb{E}_{\mathbf{W}}[\sum_{1\leq i,j\leq n}(W_{i}-1)(W_{j}-1)a_{ij}]$, and the last step holds due to the well-known relationship between operator norm and Frobenius norm $\|A\|_{\mathrm{op}}\leq\|A\|_{\mathrm{F}}$ (Golub and van Loan,, 2013, Eq. 2.3.7). Moreover, we can compute $\mu$ using explicit expressions for the joint central moments of multinomial random variables (Ouimet,, 2020, Theorem 2) as follows
@@
implies that $\Pr\nolimits_{\mathbf{W}}(|\mathbf{W}^{\top}A\mathbf{W}-\mu|\geq t)\leq\delta$, which would complete the proof by setting $C=\max(C_{1}^{1/2}C_{2}^{1/2},C_{1})$. To see this, we note that the the following are equivalent
$\delta\;\geq\;2\exp\left(-\frac{1}{C_{1}}\min\left(\frac{n^{4}t^{2}}{C_{2}\|A\|_{\mathrm{F}}^{2}},\;\frac{n^{2}t}{\|A\|_{\mathrm{F}}}\right)\right)$
@@
To bound $q^{2}_{\infty,1-\alpha/2}(\mathbb{X}_{n,\mathbf{z}})$, we use a similar argument as in Schrab et al., (2023, Appendix E4) by bounding the exceedance probability of $D^{2}_{\mathbf{W}}(\mathbb{X}_{n,\mathbf{z}})$. Notably, their proof cannot be applied directly here, because they use a wild bootstrap (Leucht and Neumann,, 2013) that corresponds to replacing $\mathbf{W}_{i}-1$ with independent Rademacher weights, whereas in our case $W_{i}$ are correlated and multinomially distributed. To proceed, we first note that we can write $D^{2}_{\mathbf{W}}(\mathbb{X}_{n,\mathbf{z}})$ as a quadratic form as $D_{\mathbf{W}}^{2}(\mathbb{X}_{n,\mathbf{z}})=n^{-2}\mathbf{W}^{\top}A\mathbf{W}$, where $A=(a_{ij})_{ij}$ with $a_{ij}\coloneqq u_{p}(\mathbf{X}_{i},\mathbf{X}_{j})$. We can then use the Hanson-Wright inequality stated in Lemma˜10 applied with $\delta=\alpha/2$ to show that there exist positive constant $C$ such that, for any almost sure realizations $\mathbb{X}_{n,\mathbf{z}}$,
@@
We now bound $\|A\|_{\mathrm{F}}^{2}$. We first show that $\Sigma_{P}^{2}\coloneqq\mathbb{E}_{\mathbf{X},\mathbf{X}^{\prime}\sim P}[u_{p}(\mathbf{X},\mathbf{X}^{\prime})^{2}]$ is finite, where $\mathbf{X},\mathbf{X}^{\prime}\sim P$ are independent. Since $u_{p}$ is a reproducing kernel (Barp et al.,, 2024, Theorem 1), we can use the reproducing property and the Cauchy-Schwarz inequality to yield
@@
$\|A\|_{\mathrm{F}}^{2}\;$ $=\;\sum_{1\leq i,j\leq n}u_{p}(\mathbf{X}_{i},\mathbf{X}_{j})^{2}$

### v5

@@ 5.3 A Kernel Exponential Family Model for Density Estimation
Our next example concerns density estimation using a kernel exponential family (KEF) model (Canu and Smola, 2006). Given a reproducing kernel $k_{0}:\mathbb{R}^{d}\times\mathbb{R}^{d}\to\mathbb{R}$ and a reference density $p_{0}$ on $\mathbb{R}^{d}$, we have $p_{\eta}(\mathbf{x})\propto p_{0}(\mathbf{x})\exp(-f_{\eta}(\mathbf{x}))$, where $f_{\eta}$ lies in the RKHS $\mathcal{H}_{k_{0}}$ associated with $k_{0}$. We follow the setting in Matsubara et al. (2022) which uses a finite-basis approximation of the RKHS function so that $f_{\eta}(\mathbf{x})=\sum_{l=1}^{25}\eta_{l}\phi_{l}(\mathbf{x})$, where $\eta=(\eta_{1},\ldots,\eta_{25})^{\top}\in\mathbb{R}^{25}$ and $\phi_{l}(\mathbf{x})=(x^{l}/\sqrt{l!})\exp(-x^{2}/2)$ for $l=1,2,\ldots,25$.
@@
We are interested in testing the goodness-of-fit of the KEF model under data contamination. Suppose that the practitioners believe the observed data are subject to contamination, so that a robust estimator for $\eta$ is used. Robust parameter estimation for exponential family models was studied in Barp et al. (2019), which proposed a minimum-distance estimator using KSD with tilted kernels. With a suitable weighting function, the resulting estimator is bias-robust (Barp et al., 2019, Proposition 7) and thus not susceptible to contamination. However, how to test the goodness-of-fit of the estimated model in this setting remains an open question. The standard IMQ-KSD test is not suitable since it is not robust and can thus falsely reject a good model due to contamination. Using a tilted kernel is also not enough since it is not quantitatively robust and hence can suffer from uncontrolled Type-I error. We therefore propose to use our robust-KSD test for this task. One challenge with the KEF model is that its density does not have a closed form, and although sampling from $P$ is feasible in this low-dimensional example ($d=1$), it would become challenging in high dimensions, making MMD-based tests not well-suited for this model. Our robust-KSD test, however, does not suffer from this limitation.
@@
We use the data set as Matsubara et al. (2022); Key et al. (2025), which is a $1$-dimensional data set of $82$ galaxy velocities (Postman et al., 1986; Roeder, 1990). To avoid using the same data for model training and testing, we randomly split the data into equal halves, each containing $n_{\mathrm{data}}=41$ data points. To each half, we then add $n_{\mathrm{ol}}$ independent draws from $R=\mathcal{N}(z,0.1^{2})$ for some value of $z$ for the mean contamination, so that the resulting data set has a size of $n=n_{\mathrm{data}}+n_{\mathrm{ol}}$ and a contamination ratio of $\epsilon=n_{\mathrm{ol}}/(n_{\mathrm{data}}+n_{\mathrm{ol}})$. We then fit the KEF model to one half of the data, and perform the tests on the other half. Both the robust minimum KSD estimation and the robust-KSD test use the same tilted kernel $k(x,y)=w(x)h_{\Lambda}(x-y)w(y)$ with $w(x)=(1+x^{2})^{-1/2}$ and a sum of IMQ kernels $h_{\Lambda}(x-y)=\sum_{\lambda^{2}\in\Lambda}(1+|x-y|^{2}/(2\lambda^{2}))^{-1/2}$ where $\Lambda=\{0.6,1,1.2\}$. The choice of the weighting function follows Matsubara et al. (2022) and the use of a sum kernel is recommended in Key et al. (2025). Direct computation shows that the Stein kernel $u_{p}$ is bounded in this case. We choose $\epsilon_{0}=0.2$ when computing $\theta$ for R-KSD.
@@ 5.4 Limitations for Multimodal Models
We investigate the performance of the robust test when the model has multimodality and the data are drawn from the same model with misspecified mixing ratios, thus revealing a limitation of our robust-KSD test. It is well-known that the standard KSD test performs poorly with mixture models with well-separated modes (Liu et al., 2023). This is intuitively due to the myopia of the score function, making the test blind to disparities in the mixing ratios, and it is a general issue for KSD and other score-based discrepancies (Wenliang and Kanagawa, 2020; Zhang et al., 2022). We demonstrate that our robust test can also suffer from this limitation, which is made even more prominent due to the relaxation of the point null to a KSD-ball. We consider a 2-dimensional mixture of Gaussian model $P=\sum_{j=1}^{5}\pi_{j}\mathcal{N}(\gamma\mu_{j},I_{2})$, where $\pi_{j}\propto u_{j}$ with $u_{1},\ldots,u_{5}$ drawn independently from $\mathrm{Uniform}(0,1)$, each mean $\mu_{j}\in\mathbb{R}^{d}$ has entries drawn independently from $\mathrm{Uniform}(-2,2)$, and $\gamma>0$ is a scaling factor controlling the mode separation. The data-generating distribution $Q$ has a different set of mixture ratios that are randomly generated. For R-KSD, the uncertainty radius is chosen to be $\theta=\epsilon_{0}\tau_{\infty}^{1/2}$ with different values of $\epsilon_{0}$, and we label the corresponding result as R-KSD-$\epsilon_{0}$.
@@ 6 Conclusion and Discussion
This paper studied the robustness of the kernel GOF test of Liu et al. (2016); Chwialkowski et al. (2016). We showed that existing KSD-based tests lack qualitative robustness when stationary kernels are used, but achieve it when employing tilted kernels inspired by robust estimation methods (Barp et al., 2019; Matsubara et al., 2022). Since qualitative robustness alone does not ensure calibration under fixed contamination levels, we proposed a novel robust KSD test that provably controls the Type-I error when the data-generating distribution lies in a KSD-ball around the reference model, while consistently rejecting alternatives outside of it. We then discussed how to select the radius of this ball under different contamination models, namely Huber’s contamination model and the density-band model. Empirical results on synthetic and real data sets demonstrate that our robust GOF test is well-calibrated and has non-trivial power. We conclude by discussing promising directions for future work.
Firstly, we established consistency of the proposed robust KSD test against all alternative distributions outside the null set, i.e., $H^{\mathrm{C}}_{1}:Q\not\in\mathcal{B}^{\mathrm{KSD}}(P;\theta)$. In the broader robust testing literature, however, it is also common to restrict the alternative set to a neighborhood centered at a specific alternative $P^{\prime}\neq P$ and study minimax properties against such hypothesis (Huber, 1965; Levy, 2008; Gül and Zoubir, 2016; Sun and Zou, 2023). Understanding whether our robust KSD test achieves minimax optimality in this setting is an interesting future work.
Secondly, as discussed in Section˜2.3, our proposed robust KSD test does not require the Stein kernel $u_{p}$ to be bounded. A bounded Stein kernel is convenient as the resulting KSD-ball $\mathcal{B}^{\mathrm{KSD}}(P;\theta)$ contains many contamination models of interest, such as Huber’s model with an arbitrary noise distribution. However, this could be too strict in some applications where outliers cannot take an arbitrary value, e.g., when the domain is bounded or the data acquisition procedure is highly regulated to prevent extreme outliers. In these cases, even with an unbounded kernel, the KSD-ball could still capture all contamination models of interest. Unbounded Stein kernels possess appealing theoretical properties such as convergence-control of moments (Kanagawa et al., 2022). Exploring whether such kernels can lead to more powerful robust-KSD tests is another worthwhile future research direction.
Acknowledgments and Disclosure of Funding
@@
Contents
@@ Appendix A Proofs of Theoretical Results
This section holds proofs of the theoretical results. Throughout the appendix, given i.i.d. random variables $\mathbf{X}_{1},\ldots,\mathbf{X}_{n}\sim Q\in\mathcal{P}(\mathbb{R}^{d})$ and an integrable function $f:(\mathbb{R}^{d})^{n}\to\mathbb{R}$, we will write for brevity $\mathbb{E}[f(\mathbf{X}_{1},\ldots,\mathbf{X}_{n})]=\int_{\mathbb{R}^{d}}\cdots\int_{\mathbb{R}^{d}}f(\mathbf{x}_{1},\ldots,\mathbf{x}_{n})Q(\mathop{}\!\mathrm{d}\mathbf{x}_{1})\cdots Q(\mathop{}\!\mathrm{d}\mathbf{x}_{n})$.
@@ A.1 Proof of Theorem˜1 and Related Preliminary Results
Assume $\sigma_{\infty}^{2}\coloneqq\sup_{\mathbf{z}\in\mathbb{R}^{d}}\mathbb{E}_{\mathbf{X}\sim P}[u_{p}(\mathbf{X},\mathbf{z})^{2}]<\infty$ and suppose Assumption˜1 holds. Then, for any $\delta>0$ and integers $0<m^{\prime}<n$, the following event holds with probability at least $1-\delta$:
@@
Suppose Assumption˜1 holds. Assume (i) $\sigma_{\infty}^{2}\coloneqq\sup_{\mathbf{z}\in\mathbb{R}^{d}}\mathbb{E}_{\mathbf{X}\sim P}[u_{p}(\mathbf{X},\mathbf{z})^{2}]<\infty$ and (ii) $\xi_{P}^{2}\coloneqq\mathbb{E}_{\mathbf{X}\sim P}[u_{p}(\mathbf{X},\mathbf{X})^{2}]<\infty$. Then there exist constants $C,C_{P,1}$, and $C_{P,2}$ depending only on $P$ such that, for any $\delta>0$ and integers $0<m^{\prime}<n$, the following event holds with probability at least $1-\delta$:
@@
There exists a constant $C>0$ such that the following holds. Suppose Assumption˜1 holds. Also assume the following integrability conditions
@@ A.1.1 Proof of Lemma˜7
Proof The test statistic can be decomposed in three terms: a term with uncorrupted samples, a term with corrupted samples, and a term for interactions between these two groups:
@@
Bounding the second term in (10) requires bounding $S_{m,\mathbf{z}}$. Because by assumption $\mathbf{X}_{i}=\mathbf{X}_{i}^{\ast}\sim P$ are i.i.d. for $i=1,\ldots,m$, the term $S_{m,\mathbf{z}}$ is a sum of i.i.d. random variables $u_{p}(\mathbf{X}_{i}^{\ast},\mathbf{z})$. Moreover, $u_{p}(\mathbf{X}_{i}^{\ast},\mathbf{z})$ are zero-mean, because the proof of Gorham and Mackey (2017, Proposition 1) shows that $\mathbb{E}_{\mathbf{X}^{\ast}\sim P}[u_{p}(\mathbf{X}^{\ast},\cdot)]=0$ whenever $\mathbb{E}_{\mathbf{X}^{\ast}\sim P}[\|\mathbf{s}_{p}(\mathbf{X}^{\ast})\|_{2}]<\infty$, which holds under Assumption˜1. We can therefore compute the variance of $S_{m,\mathbf{z}}$ as
@@ A.1.2 Proof of Lemma˜8
To prove this lemma, we will need the following concentration bound for quadratic forms of multinomial random vectors. This is a special instance of the Hanson-Wright inequality due to Adamczak (2015) when the dependent random vectors follow multinomial distributions.
Lemma 10 (Hanson-Wright inequality; Adamczak 2015, Theorem 2.5).
@@
where $\mu\coloneqq-n^{-1}\sum_{1\leq i,j\leq n}a_{ij}+\sum_{1\leq i\leq n}a_{ii}$, and $\|A\|_{\mathrm{F}}=\big(\sum_{1\leq i,j\leq n}a_{ij}^{2}\big)^{1/2}$ is the Frobenius norm.
Proof of Lemma˜10 Adamczak (2015, Remark 2.2) shows that multinomial random vectors satisfy the convex concentration inequality (Adamczak, 2015, Definition 2.2). Therefore, we can apply Adamczak (2015, Theorem 2.5) to conclude that there exists constants $C_{1},C_{2}$ such that, for any $n\times n$ matrix $A$ and any $t>0$,
@@
where $\|A\|_{\mathrm{op}}=\sup_{\|\mathbf{x}\|_{2}\leq 1}\|A\mathbf{x}\|_{2}$ is the operator norm, $\mu=\mathbb{E}_{\mathbf{W}}\big[\sum_{1\leq i,j\leq n}(W_{i}-1)(W_{j}-1)a_{ij}\big]$, and the last step holds due to the well-known ordering between operator norm and Frobenius norm $\|A\|_{\mathrm{op}}\leq\|A\|_{\mathrm{F}}$ (Golub and van Loan, 2013, Eq. 2.3.7). Moreover, we can compute $\mu$ using explicit expressions for the joint central moments of multinomial random variables (Ouimet, 2020, Theorem 2) as follows
@@
implies that $\Pr\nolimits_{\mathbf{W}}(|\mathbf{W}^{\top}A\mathbf{W}-\mu|\geq t)\leq\delta$, which would complete the proof by setting $C=\max\big(C_{1}^{1/2}C_{2}^{1/2},C_{1}\big)$. To see this, we note that the the following are equivalent
$\qquad\quad\;\delta\;\geq\;2\exp\left(-\frac{1}{C_{1}}\min\left(\frac{n^{4}t^{2}}{C_{2}\|A\|_{\mathrm{F}}^{2}},\;\frac{n^{2}t}{\|A\|_{\mathrm{F}}}\right)\right)$
@@
To bound $q^{2}_{\infty,1-\alpha/2}(\mathbb{X}_{n,\mathbf{z}})$, we use a similar argument as in Schrab et al. (2023, Appendix E4) by bounding the exceedance probability of $D^{2}_{\mathbf{W}}(\mathbb{X}_{n,\mathbf{z}})$. Notably, their proof cannot be applied directly here, because they use a wild bootstrap (Leucht and Neumann, 2013) that corresponds to replacing $\mathbf{W}_{i}-1$ with independent Rademacher weights, whereas in our case $W_{i}$ are correlated and multinomially distributed. To proceed, we first note that we can write $D^{2}_{\mathbf{W}}(\mathbb{X}_{n,\mathbf{z}})$ as a quadratic form as $D_{\mathbf{W}}^{2}(\mathbb{X}_{n,\mathbf{z}})=n^{-2}\mathbf{W}^{\top}A\mathbf{W}$, where $A=(a_{ij})_{ij}$ with $a_{ij}\coloneqq u_{p}(\mathbf{X}_{i},\mathbf{X}_{j})$. We can then use the Hanson-Wright inequality stated in Lemma˜10 applied with $\delta=\alpha/2$ to show that there exist positive constant $C$ such that, for any almost sure realizations $\mathbb{X}_{n,\mathbf{z}}$,
@@
We now bound $\|A\|_{\mathrm{F}}^{2}$. We first show that $\Sigma_{P}^{2}\coloneqq\mathbb{E}_{\mathbf{X},\mathbf{X}^{\prime}\sim P}[u_{p}(\mathbf{X},\mathbf{X}^{\prime})^{2}]$ is finite, where $\mathbf{X},\mathbf{X}^{\prime}\sim P$ are independent. Since $u_{p}$ is a reproducing kernel (Barp et al., 2024, Theorem 1), we can use the reproducing property and the Cauchy-Schwarz inequality to yield
@@
$\;\|A\|_{\mathrm{F}}^{2}$
$=\;\sum_{1\leq i,j\leq n}u_{p}(\mathbf{X}_{i},\mathbf{X}_{j})^{2}$